The time of flight of projectile motion is given by the expression?
1
Expert's answer
2020-12-15T11:47:57-0500

The time of flight of projectile motion is given by the expression (see https://courses.lumenlearning.com/boundless-physics/chapter/projectile-motion/):


"T = \dfrac{2u\sin\theta}{g}"

where "u" is the initial speed, "\theta" is the launch angle, and "g" is the gravitational acceleration.